You decide to invest in a portfolio consisting of 29 percent Stock A, 40 percent Stock B, and the remainder in Stock C. Based on the following information, what is the variance of your portfolio?

State of Economy Probability of State Return if State Occurs
of Economy Stock A Stock B Stock C
Recession .107 9.40% 2.80% 11.80%
Normal .651 8.70% 10.52% 16.20%
Boom .242 21.41% 24.83% 29.53%

Multiple Choice

  • .00867

  • .01115

  • .00826

  • .00768

  • .00933
Answer #1

Answer is 0.00826

Weight of Stock A = 0.29
Weight of Stock B = 0.40
Weight of Stock C = 0.31

Recession:

Expected Return = 0.29 * (-0.0940) + 0.40 * (-0.0280) + 0.31 * (-0.1180)
Expected Return = -0.0750

Normal:

Expected Return = 0.29 * 0.0870 + 0.40 * 0.1052 + 0.31 * 0.1620
Expected Return = 0.1175

Boom:

Expected Return = 0.29 * 0.2141 + 0.40 * 0.2483 + 0.31 * 0.2953
Expected Return = 0.2530

Expected Return of Portfolio = 0.1070 * (-0.0750) + 0.6510 * 0.1175 + 0.2420 * 0.2530
Expected Return of Portfolio = 0.1297

Variance of Portfolio = 0.1070 * (-0.0750 - 0.1297)^2 + 0.6510 * (0.1175 - 0.1297)^2 + 0.2420 * (0.2530 - 0.1297)^2
Variance of Portfolio = 0.00826